# Shannon Capacity of Channels with Markov Insertions, Deletions and Substitutions

### v1

@@ 4.1 The IDS channel sequence
Since all the terms of the summation are positive, we only must prove that the sum is not +$\infty$+ ∞ to prove convergence. Denote by $p_{n}=\sum_{y\in{\mathcal{Y}}^{n}}w_{y}$ the probability that the length of output is $n$. From (18),
@@
The second term is a finite constant. The first term is the entropy of distribution $p_{n}$, corresponding to the length of the output of $W$, which has finite expectation $\leq A$. The maximum possible entropy of discrete distribution over ${\mathbb{N}}\cup\left\{{0}\right\}$ with given (finite) mean is equal to the entropy of the geometric distribution, which is finite. ∎
4.2 The Markov-IDS channel sequence
@@
$W_{n}[\sigma]={\mathbf{g}_{n}}{\left({W_{n,1}[\sigma]}\right)}.$ (23)
@@
The Markov-IDS channel sequence $V[\rho]$, $V_{n}[\rho]={\mathbf{g}_{n}}{\left({V_{n,1}[\rho]}\right)}:{\mathcal{X}}^{n}\to{\overline{\mathcal{Y}}}$ is given by
@@
5 Capacity theorem for Markov-IDS channel
@@
5.1 Markov chain state distribution convergence
It is well-known (Theorem 8.9 in [19]), that if a Markov chain is aperiodic and irreducible, then
$\exists B\geq 0,0\leq P<1:\forall n\in{\mathbb{N}},\forall i,j\in{\mathcal{S}}:|G^{n}_{i,j}-\pi_{j}|\leq BP^{n}=B^{\prime}e^{-Cn}\stackrel{{\scriptstyle\triangle}}{{=}}\delta_{n}$ (26)
@@ 5.1 Markov chain state distribution convergence
$\varepsilon_{n}\leq De^{-Cn},C=\operatorname{const},D=\operatorname{const},$
@@ 5.2 Main Theorems
where term $f(m+n)$ is such that series $\sum_{n=1}^{\infty}f(n)/n^{2}$ converges to a finite number, then the limit $\lim_{n\to\infty}u(n)/n$ exists.
@@
Denote by $V_{n_{1}|n_{2}|...|n_{t}}[\rho]:{\mathcal{X}}^{n}\rightsquigarrow{\overline{\mathcal{Y}}}^{t}$, $n=\sum_{i}n_{i}$ a channel, which works as a channel $V_{n,1}[\rho]$ combined with the merging operation. The merging operation concatenates the outputs inside $i$-th output sub-block for each $i$, where the corresponding $i$-th input sub-block has length $n_{i}$. So, the receiver knows the output, corresponding to each $n_{i}$-block, but does not know the exact output for each input symbol. Also, we write ${\widetilde{n}}_{i}$ instead of $n_{i}$ to denote the fact that the output, corresponding to the $i$-th input block is erased, i.e., substituted by the zero-length vector. Consider the case of $\rho=\pi$, i.e., the stationary distribution of the Markov chain (later we will show that the starting distribution does not influence the i-capacity). Then, the following sequence of inequalities holds  [^4] Here transitions $1$ and $2$ correspond to providing the receiver with the additional information. In transition $3$, let us assume the same input distribution $p\in{\mathbb{D}}[{\mathcal{X}}^{m+n}]$ for both $V_{m|l|n-l}[\pi]$ and $V_{m|{\widetilde{l}}|n-l}[\pi]$ and prove that ${\mathbf{I}}(V_{m|l|n-l}[\pi],p)\leq{\mathbf{I}}(V_{m|{\widetilde{l}}|n-l}[\pi],p)+l{\overline{H}}$. If it holds for any $p$, then we establish that the transition $3$ is also valid. Denote by $X_{1},X_{2},X_{3}$ the random variables corresponding to three input blocks of $m,l$ and $(n-l)$ symbols (they are the same for both channels since we take the same input distribution), and denote the random variables corresponding to the outputs by $Y_{1},Y_{2},Y_{3}$. Then, the amount of information about $X_{1}^{3}$ that we have lost by erasing $Y_{2}$ is given by ${\mathbf{I}}(V_{m|l|n-l}[\pi],p)-{\mathbf{I}}(V_{m|{\widetilde{l}}|n-l}[\pi],p)$ $=I(Y_{1}^{3};X_{1}^{3})-I(Y_{1},Y_{3};X_{1}^{3})=I(Y_{2};X_{1}^{3}|Y_{1},Y_{3})\leq H(Y_{2}|Y_{1},Y_{3})\leq H(Y_{2}),$ where $H(Y_{2})$ is the unconditional entropy of output corresponding to the second block; vector $Y_{2}$ consists of $l$ outputs $Y_{2}^{(i)}$ of component channels, glued together, $i=1\dots l$. So, $H(Y_{2})\leq H(Y_{2}^{(1)},Y_{2}^{(2)},\dots,Y_{2}^{(l)})\leq\sum_{i}H(Y_{2}^{(i)})$. Each $Y_{2}^{(i)}$ is an unconditioned output of a state channel. It can be expressed as $H(Y_{2}^{(i)})={\mathbf{E}}_{\sigma_{i},x_{i}}{\left[{H(Y_{2}^{(i)}|x_{i},\sigma_{i})}\right]}\leq\max_{\sigma,x}{\mathbf{H}}{\left[{W[\sigma](y|x)}\right]}=\overline{H},$ where distribution of $\sigma_{i}$ is given by initial state distribution $\rho$ and Markov chain transition probabilities, and distribution of $x_{i}$ is given by marginalization of distribution $p$ over all coordinates except $(m+i)$-th. The above implies that $H(Y_{2})\leq l{\overline{H}}$ and the inequality $3$ holds. In transition $4$ after block of $(n-l)$ symbols we further transmit block of $l$ symbols. This cannot decrease the mutual information, since we can just ignore last block of input and output. In transition $5$ the output of $(n-l)$-block is merged with the output of the $l$-block. We can additionally tell the receiver the exact values of output for last $l$-block, and then the receiver can know the border between the two outputs. By the same consideration as in inequality $3$, the conditional entropy about the last $l$ glued outputs, which we have partially lost by gluing, is not greater than unconditional entropy of $l$ unglued outputs, which is not greater than $l\overline{H}$ bits. This holds for each input distribution, so this holds for maxima over input distributions. :
${\mathbf{I}}(V_{m+n}[\pi])$ $\stackrel{{\scriptstyle 1}}{{\leq}}{\mathbf{I}}(V_{m|n}[\pi])\stackrel{{\scriptstyle 2}}{{\leq}}{\mathbf{I}}(V_{m|l|n-l}[\pi])\stackrel{{\scriptstyle 3}}{{\leq}}{\mathbf{I}}(V_{m|{\widetilde{l}}|n-l}[\pi])+l{\overline{H}}\stackrel{{\scriptstyle 4}}{{\leq}}{\mathbf{I}}(V_{m|{\widetilde{l}}|n-l|l}[\pi])+l{\overline{H}}$
$\stackrel{{\scriptstyle 5}}{{\leq}}{\mathbf{I}}(V_{m|{\widetilde{l}}|n}[\pi])+2l{\overline{H}}$ (31)
@@
where $\overline{H}=\max_{\sigma\in{\mathcal{S}},x\in{\mathcal{X}}}{\mathbf{H}}{\left[{W[\sigma](y|x)}\right]}$ is the highest output entropy of a component channel (see Proposition 6).
@@
The channel which does not output the final state and the channel which outputs the final state are related via
@@
Now we can define the transition probabilities:
$V_{m|{\widetilde{l}}|n}[\pi](y_{1}^{2}|x_{1}^{3})\stackrel{{\scriptstyle\triangle}}{{=}}\sum_{\begin{subarray}{c}z_{1}\in\mathbf{g}^{-1}_{m}(y_{1})\\ z_{2}\in\mathbf{g}^{-1}_{n}(y_{2})\end{subarray}}\sum_{\sigma_{1}^{3}\in{\mathcal{S}}^{3}}\pi_{\sigma_{1}}W_{m}[\sigma_{1}\shortrightarrow\sigma_{2}]{\left({z_{1}|x_{1}}\right)}G^{l}_{\sigma_{2},\sigma_{3}}W_{n}[\sigma_{3}](z_{2}|x_{3})$ (36)
@@
$\stackrel{{\scriptstyle 2}}{{\leq}}(1+\varepsilon_{l})\sum_{\sigma_{1},z_{1}}\pi_{\sigma_{1}}W_{m}[\sigma_{1}]{\left({z_{1}|x_{1}}\right)}\sum_{\sigma_{3},z_{2}}\pi_{\sigma_{3}}W_{n}[\sigma_{3}](z_{2}|x_{3})=(1+\varepsilon_{l})V_{m}[\pi](y_{1}|x_{1})V_{n}[\pi](y_{2}|x_{3})$ (37)
@@
$\leq{\mathbf{I}}(U^{\prime},{\widetilde{p}})+\varepsilon_{l}(m+n)\overline{H}+\varepsilon_{l}(m+n)\overline{H}+\Gamma_{l},$ (38)
where $\Gamma_{l}=(1-\varepsilon_{l})\log_{2}(1+\varepsilon_{l})-(1+\varepsilon_{l})\log_{2}(1-\varepsilon_{l})$. The transition in $\stackrel{{\scriptstyle 1}}{{\leq}}$ is done as follows. Each $\log$ in the first term adds a negative amount, so the multiplicative term before the first logarithm is lower-bounded. The logarithm itself is an increasing function, so its argument is upper-bounded. Note that the upper bound might turn the value of logarithm to a positive value. Then the whole term under $\sum$ sign is upper-bounded, because the original term was non-positive. The second term is taken with -- sign, so all bounds are the opposite.
Bringing together (31), (38), for $u_{n}={\mathbf{I}}(V_{n}[\pi])$ one obtains
$u_{m+n}\leq u_{m}+u_{n}+2\varepsilon_{l}(m+n)\overline{H}+2l\overline{H}+\Gamma_{l}$
Let $l=\lfloor{{\sqrt{m+n}}\rfloor}$ and denote by $f(m+n)=2\varepsilon_{l}\overline{H}+2l\overline{H}+\Gamma_{l}$. Then,
$0\leq f(n)\leq 2\varepsilon_{2\sqrt{2n}}+2\sqrt{2n}\cdot\overline{H}+\Gamma_{\lfloor{{\sqrt{2n}}\rfloor}}$
Obviously, $\varepsilon_{l}\to 0$ and $\Gamma_{l}\to 0$ with $l\to\infty$. Thus, $f(n)\in O(\sqrt{n})$, so the series $\sum_{n}f(n)/n^{2}$ converges. This implies the existence of $\lim_{n\to\infty}u_{n}/n={\mathbf{I}}(V[\pi])$.
We can show that ${\mathbf{I}}(V[\rho])={\mathbf{I}}(V[\pi])$ by the same trick with skipping $\lfloor{{\sqrt{n}}\rfloor}$ symbols at the beginning. ∎
@@
Theorem 2 (C-capacity equals i-capacity).
For any $\rho$, the coding capacity of $V[\rho]$ achieves the mutual information capacity of $V[\rho]$, and thus does not depend on $\rho$:
${\mathbf{C}}(V[\rho])={\mathbf{C}}(V)={\mathbf{I}}(V)$ (39)
@@
The scheme of the proof is presented in Fig. 1.
@@
Figure 1: Scheme of the proof of Theorem 2
@@
Consider channel sequence ${\overline{V}}_{n}[\rho]=V_{l_{n}|m_{n}|l_{n}|...|m_{n}}[\rho]:{\mathcal{X}}^{n}\rightsquigarrow{\overline{\mathcal{Y}}}^{Q_{n}}$, for which the receiver will know the output corresponding to input subvectors of lengths $l_{n}$ and $m_{n}$, where
@@
and last $r_{n}=n-q_{n}(m_{n}+l_{n})$ input symbols are ignored. Note that with $n\to\infty$, the asymptotics of these variables are
@@
Below, we will omit index ${}_{n}$ and write $l,m,q$ and $Q$ meaning that they depend on $n$ by (40).
Also, consider channel sequence ${\overline{V}}^{\prime}_{n}[\rho]$, which outputs the output of ${\overline{V}}_{n}[\rho]$ if the total length of output is $\geq 2An$, and otherwise outputs the output of $V_{n}[\rho]$. Formally, let
$\overline{\mathbf{g}}_{n}^{\circ}({\overline{y}}_{1}^{Q})=\begin{cases}{\overline{y}}_{1}^{Q}&\text{if }{\mathbf{N}}(\mathbf{g}_{Q}({\overline{y}}_{1}^{Q}))\leq 2An\\ \mathbf{g}_{Q}({\overline{y}}_{1}^{Q})&\text{if }{\mathbf{N}}(\mathbf{g}_{Q}({\overline{y}}_{1}^{Q}))>2An\end{cases}$ (41)
and then channel sequence ${\overline{V}}^{\prime}[\rho]$ is defined by ${\overline{V}}_{n}^{\prime}[\rho]=\overline{\mathbf{g}}_{n}^{\circ}({\overline{V}}_{n}[\rho])$. Note that, with $n\to\infty$, function $\overline{\mathbf{g}}_{n}^{\circ}(y)$ changes the input with probability 0 , i.e., $\Pr{\left[{\overline{\mathbf{g}}_{n}^{\circ}(y)\neq y}\right]}\to 0$. Moreover, the image of short sequences does not intersect the image of long sequences, so we can use Proposition 5 and obtain
${\mathbf{I}}({\overline{V}}^{\prime}[\rho])={\mathbf{I}}({\overline{V}}[\rho]),~{}{\mathbf{C}}({\overline{V}}^{\prime}[\rho])={\mathbf{C}}({\overline{V}}[\rho])$ (42)
Now let $\overline{\mathbf{g}}$ be a function which glues the vectors of vectors that remained unglued after $\overline{\mathbf{g}}^{\circ}$:
@@
On one hand, $V_{n}[\rho]=\overline{\mathbf{g}}_{n}({\overline{V}}^{\prime}[\rho])$. On the other hand, note that
@@
Thus, using by Proposition 3 we obtain ${\mathbf{I}}({\overline{V}}^{\prime}[\rho])={\mathbf{I}}(V)$ and ${\mathbf{C}}({\overline{V}}^{\prime}[\rho])={\mathbf{C}}(V[\rho])$, and together with (42) we obtain
${\mathbf{I}}({\overline{V}}[\rho])={\mathbf{I}}(V),~{}{\mathbf{C}}({\overline{V}}[\rho])={\mathbf{C}}(V[\rho]).$ (44)
@@
Now, denote by ${\widetilde{V}}[\rho]$ the sequence of channels ${\widetilde{V}}_{n}[\rho]=V_{{\widetilde{l}}|m|{\widetilde{l}}|m|\dots}[\rho]$, where number of $l$- and $m$-blocks is $q$, and $l=l_{n}$, $m=m_{n}$, and $q=q_{n}$ are defined by (40). The difference between channels ${\widetilde{V}}_{n}[\rho]$ and ${\overline{V}}_{n}[\rho]$ is that each $l$-block is erased, so
${\mathbf{C}}({\widetilde{V}}[\rho])\leq{\mathbf{I}}({\widetilde{V}}[\rho])\leq{\mathbf{I}}({\overline{V}}[\rho]),{\mathbf{C}}({\widetilde{V}}[\rho])\leq{\mathbf{C}}({\overline{V}}[\rho])\leq{\mathbf{I}}({\overline{V}}[\rho])$ (45)
The total number of inputs, for which the outputs are erased, is $ql$. Assume some input distribution $p$ over ${\mathcal{X}}^{n}$. Then, the following random variables are defined: $X$ corresponds to the input; $Y$ is the erased output $l$-blocks; $Z$ is the output of ${\overline{V}}_{n}[\rho]$. Then,
${\mathbf{I}}({\overline{V}}_{n}[\rho],p)$ $=I(X;Y,Z)=I(X;Z)+I(X;Y|Z)={\mathbf{I}}({\widetilde{V}}_{n}[\rho],p)+I(X;Y|Z)$
$\leq{\mathbf{I}}({\widetilde{V}}_{n}[\rho],p)+H(Y)\leq{\mathbf{I}}({\widetilde{V}}_{n}[\rho],p)+lq{\overline{H}}$
Since $q\in O(\sqrt{n})$ and $l\in O(\sqrt[4]{n})$, the last term is in $o(n)$, which means that ${\mathbf{I}}({\overline{V}}[\rho])\leq{\mathbf{I}}({\widetilde{V}}[\rho])$. Recalling (45), one obtains ${\mathbf{I}}({\overline{V}}[\rho])={\mathbf{I}}({\widetilde{V}}[\rho])$.
Step 3. ${\widetilde{V}}$ is information-stable, and, thus, ${\mathbf{I}}({\widetilde{V}})={\mathbf{C}}({\widetilde{V}})$.

### v3

@@ IV-A The IDS channel sequence
Since all the terms of the summation are non-negative, we only need to prove that the sum is constrained. Denote by $p_{n}=\sum_{y\in{\mathcal{Y}}^{n}}w_{y}$ the probability that the length of output is $n$. From (23),
@@
The second term is finite. The first term is the entropy of distribution $p_{n}$, corresponding to the length of the output of $W$, which has finite expectation $\leq A$. The maximum possible entropy of discrete distribution over ${\mathbb{N}}\cup{\left\{{0}\right\}}$ with given (finite) mean is equal to the entropy of the geometric distribution, which is finite. ∎
IV-B The Markov-IDS channel sequence
@@
$W_{n}[\sigma]=\mathbf{g}_{n}(W_{n,1}[\sigma]).$ (28)
@@
The Markov-IDS channel sequence $V[\rho]$, $V_{n}[\rho]=\mathbf{g}_{n}(V_{n,1}[\rho]):{\mathcal{X}}^{n}\to{\overline{\mathcal{Y}}}$ is given by
@@
V Capacity theorem for Markov-IDS channel
@@
V-A Markov chain state distribution convergence
It is well-known (Theorem 8.9 in [21]), that if a Markov chain with a finite state space is aperiodic and irreducible, then
$\exists B\geq 0,0\leq P<1:\forall n\in{\mathbb{N}},\forall i,j\in{\mathcal{S}}:|G^{n}_{i,j}-\pi_{j}|\leq BP^{n}=B^{\prime}e^{-Cn}\stackrel{{\scriptstyle\triangle}}{{=}}\delta_{n},$ (31)
@@ V-A Markov chain state distribution convergence
$\varepsilon_{n}\leq De^{-Cn},C=\operatorname{const},D=\operatorname{const},$ (34)
@@ V-B Main Theorems
where term $f(m+n)$ is such that series $\sum_{n=1}^{\infty}f(n)/n^{2}$ converges, then the limit $\lim_{n\to\infty}u(n)/n$ exists.
@@
Denote by $V_{n_{1}|n_{2}|...|n_{t}}[\rho]:{\mathcal{X}}^{n}\rightsquigarrow{\overline{\mathcal{Y}}}^{t}$, $n=\sum_{i}n_{i}$, a channel, which works as a channel $V_{n,1}[\rho]$ combined with the merging operation. The merging operation concatenates the outputs inside $i$-th output sub-block for each $i$, where the corresponding $i$-th input sub-block has length $n_{i}$. So, the receiver knows the output, corresponding to each $n_{i}$-block, but does not know the exact output for each input symbol. Also, we write ${\widetilde{n}}_{i}$ instead of $n_{i}$ to denote the fact that the output, corresponding to the $i$-th input block is erased, i.e., substituted by the zero-length vector. Consider the case of $\rho=\pi$, i.e., the initial state distribution is the stationary distribution of the Markov chain (later we will show that the initial state distribution does not influence the i-capacity). Then, the following sequence of inequalities holds:
${\mathbf{I}}(V_{m+n}[\pi])$ $\stackrel{{\scriptstyle 1}}{{\leq}}{\mathbf{I}}(V_{m|n}[\pi])\stackrel{{\scriptstyle 2}}{{\leq}}{\mathbf{I}}(V_{m|l|n-l}[\pi])\stackrel{{\scriptstyle 3}}{{\leq}}{\mathbf{I}}(V_{m|{\widetilde{l}}|n-l}[\pi])+l{\overline{H}}\stackrel{{\scriptstyle 4}}{{\leq}}{\mathbf{I}}(V_{m|{\widetilde{l}}|n-l|l}[\pi])+l{\overline{H}}\stackrel{{\scriptstyle 5}}{{\leq}}{\mathbf{I}}(V_{m|{\widetilde{l}}|n}[\pi])+2l{\overline{H}}$ (37)
@@
where $\overline{H}=\max_{\sigma\in{\mathcal{S}},x\in{\mathcal{X}}}{\mathbf{H}}{\left[{W[\sigma](y|x)}\right]}$ is the highest output entropy of a component channel (see Proposition 7). Here, transitions $1$ and $2$ correspond to providing the receiver with additional information.
In transition $3$, let us assume the same input distribution $p\in{\mathbb{D}}[{\mathcal{X}}^{m+n}]$ for both $V_{m|l|n-l}[\pi]$ and $V_{m|{\widetilde{l}}|n-l}[\pi]$ and prove that ${\mathbf{I}}(V_{m|l|n-l}[\pi],p)\leq{\mathbf{I}}(V_{m|{\widetilde{l}}|n-l}[\pi],p)+l{\overline{H}}$. Since this holds for any $p$, then we establish that the transition $3$ is also valid. Denote by $X_{1},X_{2},X_{3}$ the random variables corresponding to three input blocks of $m,l$ and $(n-l)$ symbols (they are the same for both channels since we take the same input distribution), and denote the random variables corresponding to the outputs by $Y_{1},Y_{2},Y_{3}$. Then, the amount of information about $X_{1}^{3}$ that we have lost by erasing $Y_{2}$ is given by
${\mathbf{I}}(V_{m|l|n-l}[\pi],p)-{\mathbf{I}}(V_{m|{\widetilde{l}}|n-l}[\pi],p)=I(Y_{1}^{3};X_{1}^{3})-I(Y_{1},Y_{3};X_{1}^{3})=I(Y_{2};X_{1}^{3}|Y_{1},Y_{3})\leq H(Y_{2}|Y_{1},Y_{3})\leq H(Y_{2}),$
where $H(Y_{2})$ is the unconditional entropy of output corresponding to the second block; vector $Y_{2}$ consists of $l$ outputs $Y_{2}^{(i)}$ of component channels, glued together, $i=1\dots l$. So, $H(Y_{2})\leq H(Y_{2}^{(1)},Y_{2}^{(2)},\dots,Y_{2}^{(l)})\leq\sum_{i}H(Y_{2}^{(i)})$. Each $Y_{2}^{(i)}$ is an unconditioned output of a state channel. It can be expressed as
$H(Y_{2}^{(i)})={\mathbf{E}}_{\sigma_{i},x_{i}}{\left[{H(Y_{2}^{(i)}|x_{i},\sigma_{i})}\right]}\leq\max_{\sigma,x}{\mathbf{H}}{\left[{W[\sigma](y|x)}\right]}=\overline{H},$
where distribution of $\sigma_{i}$ is given by initial state distribution $\rho$ and Markov chain transition probabilities, and distribution of $x_{i}$ is given by marginalization of distribution $p$ over all coordinates except the $(m+i)$-th coordinate. The above implies that $H(Y_{2})\leq l{\overline{H}}$ and inequality $3$ holds.
In transition $4$ after block of $(n-l)$ symbols we further transmit a block of $l$ symbols. This cannot decrease the mutual information, since we can just ignore last block of input and output.
In transition $5$ the output of $(n-l)$-block is merged with the output of the $l$-block. We can additionally tell the receiver the exact values of the output for the last $l$-block, and then the receiver can know the boundaries of the two outputs. By the same consideration as in inequality $3$, the conditional entropy about the last $l$ glued outputs, which we have partially lost by gluing, is not greater than unconditional entropy of $l$ unglued outputs, which is not greater than $l\overline{H}$ bits. This holds for any input distribution, so this holds for maxima over the input distributions.
@@
The channel which does not output the final state and the channel which outputs the final state are related via marginalization over the final state:
@@
Now we can define the transition probabilities, corresponding to transmitting an $m$ and an $n$-block, separated by an $l$-block, which is then erased:
$V_{m|{\widetilde{l}}|n}[\pi](y_{1}^{2}|x_{1}^{3})\stackrel{{\scriptstyle\triangle}}{{=}}\sum_{\begin{subarray}{c}z_{1}\in\mathbf{g}^{-1}_{m}(y_{1})\\ z_{2}\in\mathbf{g}^{-1}_{n}(y_{2})\end{subarray}}\sum_{\sigma_{1}^{3}\in{\mathcal{S}}^{3}}\pi_{\sigma_{1}}W_{m}[\sigma_{1}\shortrightarrow\sigma_{2}]{\left({z_{1}|x_{1}}\right)}G^{l}_{\sigma_{2},\sigma_{3}}W_{n}[\sigma_{3}](z_{2}|x_{3})$
@@
$\stackrel{{\scriptstyle 2}}{{\leq}}(1+\varepsilon_{l})\sum_{\sigma_{1},z_{1}}\pi_{\sigma_{1}}W_{m}[\sigma_{1}]{\left({z_{1}|x_{1}}\right)}\sum_{\sigma_{3},z_{2}}\pi_{\sigma_{3}}W_{n}[\sigma_{3}](z_{2}|x_{3})=(1+\varepsilon_{l})V_{m}[\pi](y_{1}|x_{1})V_{n}[\pi](y_{2}|x_{3}).$ (40)
@@
$\leq{\mathbf{I}}(V_{m}[\pi]\cdot V_{n}[\pi],{\widetilde{p}})+\varepsilon_{l}(m+n)\overline{H}+\varepsilon_{l}(m+n)\overline{H}+\Gamma_{l},$ (41)
where $\Gamma_{l}=(1-\varepsilon_{l})\log_{2}(1+\varepsilon_{l})-(1+\varepsilon_{l})\log_{2}(1-\varepsilon_{l})$. The transition in $\stackrel{{\scriptstyle 1}}{{\leq}}$ is done as follows. Each $\log$ in the first term adds a negative amount, so the multiplicative term before the first logarithm is lower-bounded. The logarithm itself is an increasing function, so its argument is upper-bounded. Note that the upper bound might turn the value of logarithm to a positive value. Then the whole term under $\sum$ sign is upper-bounded, because the original term was non-positive. The second term is taken with a negative sign, so all bounds are the opposite.
Bringing together (37) and (41), for $u_{n}={\mathbf{I}}(V_{n}[\pi])$ one obtains
$u_{m+n}\leq u_{m}+u_{n}+2\varepsilon_{l}(m+n)\overline{H}+\Gamma_{l}+2l\overline{H}$
Let $l=\lfloor{{\sqrt{m+n}}\rfloor}$ and denote by $f(m+n)=2\varepsilon_{l}(m+n)\overline{H}+\Gamma_{l}+2l\overline{H}$. Then,
$0\leq f(n)\leq 2\varepsilon_{2\sqrt{2n}}\cdot n+\Gamma_{\lfloor{{\sqrt{2n}}\rfloor}}+2\sqrt{2n}\cdot\overline{H}$
Note that by (34), the first term tends to zero. The second term $\Gamma_{l}\to 0$ with $l\to\infty$. Thus, $f(n)\in O(\sqrt{n})$, so the series $\sum_{n}f(n)/n^{2}$ converges. This implies the existence of $\lim_{n\to\infty}u_{n}/n={\mathbf{I}}(V[\pi])$.
We can show that ${\mathbf{I}}(V[\rho])={\mathbf{I}}(V[\pi])$ by the same trick of skipping $\lfloor{{\sqrt{n}}\rfloor}$ symbols at the beginning. ∎
@@
Theorem 2 (Information stability of the Markov-IDS channel sequence).
For any $\rho$, the channel sequence $V[\rho]$ is information stable. In particular, the coding capacity ${\mathbf{C}}(V[\rho])$ achieves the mutual information capacity ${\mathbf{I}}(V)$, and thus does not depend on $\rho$:
${\mathbf{C}}(V[\rho])={\mathbf{C}}(V)={\mathbf{I}}(V).$ (42)
@@
The outline of the proof is depicted in Fig. 1.
@@
Figure 1: The outline of proof of Theorem 2
@@
Consider channel sequence ${\overline{V}}_{n}[\rho]=V_{l_{n}|m_{n}|l_{n}|...|m_{n}}[\rho]:{\mathcal{X}}^{n}\rightsquigarrow{\overline{\mathcal{Y}}}^{Q_{n}}$, for which the receiver knows the outputs corresponding to all the input subvectors of lengths $l_{n}$ and $m_{n}$, where
@@
and the last $r_{n}=n-q_{n}(m_{n}+l_{n})$ input symbols are ignored. Note that with $n\to\infty$, the asymptotics of these variables are
@@
Below, we will omit index n and write $l,m,q$ and $Q$ meaning that they depend on $n$ by (43).
Note that in the case of channel ${\overline{V}}_{n}[\rho]$ the receiver knows all the information that is here in the case of channel $V_{n}[\rho]$ plus the boundaries of $l$- and $m$-blocks. This means that channel $V_{n}[\rho]$ can be represented as a pipeline of channel ${\overline{V}}_{n}[\rho]$ and merging function $\mathbf{g}_{Q}$, which merges $Q$ vectors into one, i.e., $V_{n}[\rho]=\mathbf{g}_{Q}({\overline{V}}_{n}[\rho])$.
We propose to represent the merging function $\mathbf{g}_{Q}$ as a pipeline of two functions. The first one merges the output of ${\overline{V}}_{n}[\rho]$, if the total length of output is $<2An$. Formally, let
$\mathbf{g}_{n}^{\circ}({\overline{y}}_{1}^{Q})=\begin{cases}{\overline{y}}_{1}^{Q}&\text{if }{\mathbf{N}}(\mathbf{g}_{Q}({\overline{y}}_{1}^{Q}))\leq 2An\\ \mathbf{g}_{Q}({\overline{y}}_{1}^{Q})&\text{if }{\mathbf{N}}(\mathbf{g}_{Q}({\overline{y}}_{1}^{Q}))>2An\end{cases}$ (44)
Note that, with $n\to\infty$, function $\mathbf{g}_{n}^{\circ}(y)$ changes the input with probability 0 , i.e., $\Pr{\left[{\mathbf{g}_{n}^{\circ}(y)\neq y}\right]}\to 0$. Moreover, the image of short sequences does not intersect the image of long sequences, so function sequence $\mathbf{g}^{\circ}$ satisfy Proposition 6.
Now let $\overline{\mathbf{g}}$ be a sequence of functions, which glue the vectors of vectors that remained unglued after $\mathbf{g}^{\circ}$:
@@
Note that the log-size of the inverse image of this function is
@@
Thus, function sequence $\overline{\mathbf{g}}$ satisfies Proposition 4. One can see that $\mathbf{g}_{Q}=\mathbf{g}^{\circ}_{n}\circ\overline{\mathbf{g}}_{n}$, thus, $V[\rho]=\overline{\mathbf{g}}(\mathbf{g}^{\circ}({\overline{V}}[\rho]))$, and we obtain
${\mathbf{I}}({\overline{V}}[\rho])={\mathbf{I}}(V),~{}{\mathbf{I}}{\mathbf{S}}({\overline{V}}[\rho])\iff{\mathbf{I}}{\mathbf{S}}(V[\rho]).$ (46)
@@
Now, denote by ${\widetilde{V}}[\rho]$ the sequence of channels ${\widetilde{V}}_{n}[\rho]=V_{{\widetilde{l}}|m|{\widetilde{l}}|m|\dots}[\rho]$, where number of $l$- and $m$-blocks is $q$, and $l=l_{n}$, $m=m_{n}$, and $q=q_{n}$ are defined by (43).
The difference between channels ${\widetilde{V}}_{n}[\rho]$ and ${\overline{V}}_{n}[\rho]$ is that each $l$-block is erased. Similarly to the previous step, transition from ${\overline{V}}_{n}[\rho]$ to ${\widetilde{V}}_{n}[\rho]$, i.e., erasing of the $l$-blocks, can be done by two steps via pipeline of two functions ${\mathbf{f}}^{\circ}_{n}$ and $\overline{{\mathbf{f}}}_{n}$, such that ${\widetilde{V}}_{n}[\rho]=\overline{{\mathbf{f}}}_{n}({\mathbf{f}}^{\circ}_{n}({\widetilde{V}}_{n}[\rho]))$.
Let function ${\mathbf{f}}^{\circ}_{n}$ substitute all $l$-blocks with special erasure symbol $\iota\notin{\mathcal{Y}}$, if their total length is greater than $2Aql$. Similarly to $\mathbf{g}^{\circ}$, function sequence ${\mathbf{f}}^{\circ}$ satisfies conditions for Proposition 6.
Let function $\overline{{\mathbf{f}}}_{n}$ erase all $l$-blocks after applying function ${\mathbf{f}}^{\circ}_{n}$. Then, the number $\Phi_{n}$ of possible inputs, corresponding to the same output, or, alternatively, the number of possible values of erased $l$-blocks, is upper-bounded by $\sum_{i=0}^{\lfloor{{2Aql}\rfloor}}|{\mathcal{Y}}\cup{\left\{{\iota}\right\}}|^{i}\leq(|{\mathcal{Y}}|+1)^{2Aql+1}$. Obviously, $A$ and $|{\mathcal{Y}}|$ are constants, and $ql\in O(n^{3/4})$, thus, $\log_{2}\Phi_{n}\in O(n^{3/4}\log_{2}n)\subset o(n)$, which means that function sequence $\overline{{\mathbf{f}}}$ satisfies Proposition 4.
The above considerations imply
${\mathbf{I}}({\widetilde{V}}[\rho])={\mathbf{I}}({\overline{V}}[\rho]),{\mathbf{I}}{\mathbf{S}}({\widetilde{V}}[\rho])\iff{\mathbf{I}}{\mathbf{S}}({\overline{V}}[\rho]).$ (47)
Step 3. Information stability of ${\widetilde{V}}$.